# An Empirical Study of Training ID-Agnostic Multi-modal Sequential Recommenders

### v2

@@ 5.4. Overall Performance Comparisons
  • Our experiments show that MoRec and our proposed method outperform VQRec [^8] We observe VQRec do not perform well under our experimental setting, which might be attributed to the following reasons: non-end-to-end training manner (Yuan et al., 2023), lack of multimodal information., indicating that end-to-end learning is more effective than traditional pre-extraction. This is likely due to the fact that end-to-end learning allows the item encoder to better capture the semantic information of modalities (Yuan et al., 2023).
6. Inherit Advances in Multi-modal Pre-training Community (Q2)
MMSR employs SR as a subsequent task in multimodal learning, prompting us to examine if advancements in the multi-modal pre-training community, i.e., Vision-and-Language Pre-training (VLP), can enhance MMSR. In particular, we delve into two fundamental questions frequently addressed in the VLP community: (1) the architecture of VLP model (Singh et al., 2022; Kim et al., 2021), and (2) the visual representation within VLP model (Kim et al., 2021).
Table 6. Comparison of MMSR with various VLP models of different architectures. (a): two-stream & shallow fusion. (b): single stream & deep fusion. (c): two-stream & deep fusion.
@@ 6. Inherit Advances in Multi-modal Pre-training Community (Q2)
Table 7. Comparison of vision representations for MMSR.
@@
6.1. Impact of Various VLP Architectures
Our exploration initiates with a meticulous examination of diverse VLP architectures. To be specific, these can be broadly classified into three emblematic paradigms: (1) Two-stream and shallow fusion, as exemplified by models such as CLIP (Radford et al., 2021) and GroupViT (Xu et al., 2022), (2) Single-stream and deep fusion, represented by the likes of ViLT (Kim et al., 2021) and VisualBERT (Li et al., 2019), and (3) Two-stream and deep fusion, encapsulated by models like FLAVA (Singh et al., 2022).
For architectures such as CLIP (Radford et al., 2021) and GroupViT (Xu et al., 2022) that lack a fusion module, we elected to employ the widely-adopted concatenation technique to glean cross-modal representations. For all transformer-based encoders, we limited our fine-tuning to the two top-layer transform blocks, thereby adhering to our prior experiments. In the case of FLAVA, considering its substantially deep fusion module, we only fine-tune its final fusion block.
Results. In Table 6, we can see that MMSR performs worst in the CLIP & GroupViT architecture. By analyzing in conjunction with the other two VLPs, we hypothesize it is the shallow fusion that deteriorates the performance. We also notice that FLAVA drives MMSR to completely surpass other VLPs, which is consistent with its excellent performance in the multi-modal fields. Note, the best-performing MMSR in RQ(1) has a similar architecture to FLAVA, which may imply that the success of PTM in other domains can be transferred to the recommendation task.
6.2. Impact of Various Vision Representation
We carry out further exploration by evolving vision representation in VLPs (see Section 3 for details). We select VisualBERT (Li et al., 2019) as our foundational model due to its iconic status within the domain of VLPs. Most subsequent strides of vision representation are indeed enhancements built upon it. To validate the MMSR, we have incorporated an assortment of vision representations, including region-based (Faster-RCNN (Faster, 2015)), grid-based (ResNet50 (He et al., 2016)), patch-Embedding (CLIP-ViT (Radford et al., 2021), excluding 12 transformers) and patch-Vision Transformer (CLIP-ViT, including 12 transformers). These representations typify a significant trajectory in the evolution of VLPs, thereby prompting our curiosity to investigate their adaptability to MMSR.
Results. As is shown in Table 7, all MMSRs with various vision representations can exceed IDSR, and patch-ViT representation is the best, which is in line with the development process in the field of VLPs. Region-based and Patch-Emd representations underperform, whereas MMSR with Grid-based and Patch-ViT demonstrate notable improvements over IDSR. This disparity might be attributed to the strong augmentation of the latter two specific vision encoders, namely ResNet50 (He et al., 2016) and CLIP-ViT (Radford et al., 2021). Note that in Section 5.1, ResNet50 (grid-based) performs poorly in the previous setting. However, it achieves good results here, suggesting that deep fusion may resolve instability caused by inconsistent representations.
7. Cold-start and Transferability (Q3)
7.1. MMSR in Cold-start Scenario
Table 8. Performance comparison (%) under cold-start setting.
@@ 7.1. MMSR in Cold-start Scenario
NDCG@10 0.1381 4.8547 4.3779 5.4046
Results. Table 8 reveals that MMSR, MMSR-T, and MMSR-V all notably outperform IDSR in predicting cold items. This superior performance stems from the fact that IDSR only derives representations from behavioral patterns, but MMSR incorporates original modal features, the advantage of which becomes more apparent for cold start. Interestingly, while MMSR-T and MMSR-V have comparable performances in standard situations (as is shown in Table 4), MMSR-T tends to be more adept at handling cold items. This suggests that text-based representations might have a stronger generalizing capability for recommendation tasks than visual ones.
7.2. Results of Transfer Learning
@@
Scaling the source domain. As shown in Table 9, we choose the HM-100K, -200K, and -500K datasets as the source domain and the e-commerce Amazon-10K (He and McAuley, 2016; McAuley et al., 2015) dataset as the target domain. As the scale of the source domain increases, the results of transferring increase compared with those from scratch. This implies that by increasing the sample size, MMSR can more effectively uncover user preferences within the modal content.
Scaling the target domain. On the contrary, we select Kwai-200K as the source dataset and scale the target Bili dataset from 10K, 20K to 50K. As shown in Table 10, the transferring improvement is reduced as the target domain expands.
Results. As is shown in Table 9 and Table 10, MMSR can perform better than training from scratch, but it does not always hold when the ratio of the source domain to the target domain is narrowed, e.g., the results of transferring from Kwai-200K to Bili-20K and -50K show no advantages. The performance drop might stem from differences in user groups between domains. Larger downstream datasets can introduce unrelated features and noise, hampering the transfer learning performances of sequential recommenders.
Table 9. Transfer learning from HM to Amazon. Source: HM-100K, HM-200K, HM-500K. Target: Amazon-10K Clothes Dataset. ‘Improv.’ is the relative improvement of the best transferring compared to training from scratch.
@@
NDCG@10 17.2302 21.7594 23.3476 22.9698 23.3558 +6.83%
Table 10. Transfer learning from Kwai to Bili. Source: Kwai-200K dataset. Target: Bili-10K, -20K, and -50K dataset. ‘Improv.’ is the relative improvement of the transferring compared to the corresponding training from scratch.
@@
8. Conclusion
Appendix A Details of Datasets
The HM and Amazon of clothes category dataset includes customer-product (user-item) interactions and customer and product metadata, e.g., category and color, descriptions, and covers, as shown in Figure 3 (left). In this paper, we only use texts and images, adding some category information as part of the description to further distinguish different items. Following (Yuan et al., 2019), we removed users with less than 5 interactions simply, and randomly selected 200K and 100K users.
To build the Bili dataset, we randomly crawl videos (less than 10 minutes in duration) of different categories from the online platform. We record the public comment information of these videos as interactions (i.e., we think the user would be interested in this video if a user comments on a video). We do not capture any user’s personal information about privacy, whose IDs are already anonymous after data masking. The video metadata such as cover and title are recorded as shown in Figure 3 (right). The Kwai dataset is collected similarly to the Bili dataset. Then we also removed users with less than 5 interactions and randomly selected about 100K and 200K users from the Bili and Kwai datasets, respectively.
Besides, for all datasets, we consider cold scenarios (different from regular scenarios). That is, we count the interactions of all items in the training set and consider those items with less than 10 occurrences as cold items (Anonymous, 2023). Then, we truncate the complete user sequence to get some sub-sequences with a cold item at the end, all of which are used to evaluate the capability of MMSR in the cold scenario.
Appendix B Details of Various SR Architectures
In this study, we undertake a comprehensive empirical analysis using three representative SR architectures to rigorously assess the efficacy of MMSR. The selected architectures encompass a broad spectrum of methodologies, specifically: the Transformer-based SASRec (Kang and McAuley, 2018), the GRU-based GRURec (Li et al., 2022), and the CNN-based NextItNet (Yuan et al., 2019). Our main objective is to explore the dynamic interrelationship between these prevalent sequential recommendation architectures and the intricacies of the multi-modal item encoder. To maintain consistent efficiency, we adopt dense next-item prediction methodologies for all these SR architectures(Kang and McAuley, 2018).
(3) $R(x)=\prod\limits_{i=0}^{l}r(x_{i}|x_{0:l-1},\theta)r(x_{0})$
where $r(x_{i}|x_{0:l-1},\theta)$ means the probability of $i$-th item conditioned on all previous item $x_{0:i-1}$.
Obviously, joint distributions are more efficient at capturing sets of all sequence relations, and transformers are more powerful than GRU and CNN basic modules. Hence, GRURec and NextItNet would be taken as degenerated versions of SASRec. Our results on the above 3 SR architectures can prove that MMSR is widely applicable in most cases. We show some specific hyperparameters for SASRec, GRU4Rec, and NextItNet. The layer number is set to 2, 4, and 2, respectively, which is from searching in $\{1,2,4,8\}$. The attention headers of SASRec are set to 2.
Appendix C Evaluations Metrics
In this paper, by employing a standard ”leave-one-out” strategy (Kang and McAuley, 2018), we split the dataset into training, validation, and test sets. For example, we selected the latest 23 items of each user sequence for all datasets here (i.e., 1-21 items for training, 2-22 for validation, and 3-23 for testing).
We evaluate all models using two Top-N ranking metrics: HR@N (Hits Ratio) and NDCG@N (Normalized Discounted Cumulative Gain) (Yuan et al., 2019), where N is set to 10 in this paper. Specifically,
(4) $HR=\frac{1}{N}\sum_{i=0}^{N}hits(i)$
$hits(i)$ predict whether the user’s $i$-th intention item is in the recommended rank list (1 if it is, and 0 otherwise).
(5) $NDCG=\frac{1}{N}\sum_{i=0}^{N}\frac{1}{log_{2}(p_{i}+1)}$
$p_{i}$ means the position of the user’s $i$-th intended item in the recommendation list. It would be close to +$\infty$+ ∞ if there is no user’s $i$-th intended item in the recommendation rank list.
Following (Krichene and Rendle, 2022), we rank the ground-truth target item by comparing it with all the other items in the item pool. Finally, we report the results of the testing set by choosing checkpoints with the best validation performance.
Appendix D Performance of MMSR (Q1)
Figure 4 shows the convergence of IDSR, MMSR-T, MMSR-V, and MMSR with the best setting in Table 2 and 4. i.e., SASRec RS architecture, RoBERTa (Liu et al., 2019) and Swin B-224/16 (Liu et al., 2021) as default text and vision encoder, respectively.
Figure 4. Convergence of IDSR, MMSR-T, MMSR-V and MMSR on HM, Bili, and Kwai datasets.
It can be seen that there is a performance gap between the validation and the test results by comparing Figure 4 and Table 4. In many validation experiments, MMSR-T and MMSR-V are not comparable to IDSR, but they are comparable or even surpassed IDSR on the test set, e.g., the HM-200K and Kwai-200K datasets. This implies the better generalization performance of the modality-based recommendation than IDSR. Especially, MMSR has always been able to greatly surpass IDSR, no matter how on the validation set and on the test set.
Appendix E MMSR with Various VLPs (Q(2))
In this paper, we introduce three typical VLP paradigms to our MMSR to see their impact on the recommendation task, i.e., CLIP (Kang et al., 2020) and GroupViT (Xu et al., 2022) (two-stream and shallow fusion), VsualBERT (Li et al., 2019) and ViLT (Kim et al., 2021) (single stream and deep fusion), and FLAVA (Singh et al., 2022) (two-stream and deep fusion), as shown in 5. We show the convergence of MMSR with the above VLPs in Figure 6.
Figure 5. Description of each module: a) CLIP-B-224/32 and GroupViT (a CLIP variant for semantic segmentation); b) VisualBERT and ViLT; c) FLAVA.
Specifically, we choose SASRec as the default RS architecture. For CLIP and GroupViT, we fine-tune two transformer blocks for the text and vision encoders and use a concatenation layer to connect both as the input of the downstream recommendation network. Since VisualBERT can not extract vision features itself, we use Faster-RCNN(Faster, 2015) with a threshold of 0.05 and select 49 regions from each image as visual input. The last 4 transformer blocks of VisualBERT are fine-tuned, which is consistent with ViLT. As for FLAVA, we fine-tune the top two transformer blocks of text and vision encoders and all fusion modules.
Figure 6. Convergence of MMSR with various VLPs of 3 different architectures.
Figure 6 demonstrates MMSR with CLIP or GroupViT performs the worst. A possible hypothesis is that shallow fusion leads to performance degradation. We note that FLAVA drives MMSR to completely outperform IDSR, which is consistent with its excellent performance in the multimodal community.
E.1. Transferring of MMSR (Q(3))
Figure 7 show the convergence of the transferring experiments across platforms of MMSR. We take HM-500K and Kwai-200K as source domain datasets, and Amazon-10K and Bili-10K as target domain datasets, respectively (see Table 9 and 10). We train IDSR for 300 epochs, and the learning rate decays linearly to 10% of the original learning rate after 270 epochs. And both MMSR-Transfer and MMSR-Scratch are only trained for 60 epochs (already converged), and the learning rate decays linearly to 10% after 30 epochs. Compared with MMSR-Scratch, MMSR-Transfer achieves surprising results, i.e., more stable, faster, and better performance.
Figure 7. Comparison of MMSR-Transfer and MMSR-Scratch. Note IDSR is trained for 300 epochs, and we only record results every 5 epochs for it.
Appendix F Bag of Tricks
@@
F.1. Full Fine-tuning vs. Top-layer Fine-tuning
In all the above experiments, we fine-tune only the top blocks for all encoders by default, which is mainly due to computational consumption. Here, we try to investigate whether all parameters need to be adapted. Table 11 shows the results of full and top-layer fine-tuning showed irregular fluctuations. Among them, top-layer fine-tuning works better in MMSR. We conclude that top-layer fine-tuning is worth considering in order to balance accuracy and efficiency in practice.
Table 11. Full vs top-layer fine-tuning.
@@
F.2. Different Learning Rates
As shown in Table 12, using the same learning rate for all modules of the model will lead to performance degradation, which may be because the pre-trained parameters already contain a certain amount of knowledge about vision and language, and over-tuning them may lead to loss of this valuable information. As shown in Table 12, the performance of MMSR would deteriorate when all modules use the same learning rate. This is probably because the pre-trained parameters in the item encoders already contain a certain amount of knowledge about vision and language, and over-tuning may lead to the loss of this knowledge. But the downstream recommendation networks and fusion modules need training from scratch, so a large learning rate would be better.
Table 12. Comparison of MMSR with different learning rates for different modules. Top: TE and VE of MMSR. Downstream: the recommendation networks and fusion modules.
@@ F.2. Different Learning Rates
F.3. Length of Texts
In our investigation, we dedicated substantial effort to analyzing the impact of varying quantities of textual information. This aspect is crucial, as the volume of textual data in a recommendation system can significantly influence its predictive accuracy. Our primary focus was on discerning how changes in the length of titles, an integral component of textual data, affected the performance of our models. As demonstrated in Table 13, our exploration yielded noteworthy results. Both MMSR-T and MMSR exhibited a positive response to increases in title length. This trend implies that an expansion of title content can contribute to enhanced model performance. Our findings underline the significance of textual volume, particularly in title length, as a determinant factor for improving the functionality of MMSR.
Table 13. The impact of various lengths of text for MMSR-T and MMSR.
@@ F.3. Length of Texts
F.4. Mask Ratio of Images
The volume of textual and visual information can greatly affect model performance MMSR. Inspired by MAE (He et al., 2022), we control the image volume by adjusting its mask ratio. Table 14 shows that as the mask ratio increases, the performance of MMSR-V rises but MMSR decreases gradually. A possible reason is that information is insufficient for MMSR-V thus it would be better to keep more visual features. While for MMSR, the information may be saturated with both linguistic and visual features being fed, and it would be more intractable to optimize since the redundant image information would be taken as noise in this case.
Table 14. The impact of the volume of image information for MMSR. Note that we use mask ratio from MAE (He et al., 2022) to control the volume of image information.

### v5

@@ 5.4 Overall Performance Comparisons
  • Our experiments show that MoRec and our proposed method outperform VQRec, indicating that end-to-end learning is more effective than traditional pre-extraction. This is likely due to the fact that end-to-end learning allows the item encoder to better capture the semantic information of modalities [8].
6 Inherit Advances in Multi-modal Pre-training Community (Q2)
MMSR employs SR as a subsequent task in multimodal learning, prompting us to examine if advancements in the multi-modal pre-training community, i.e., Vision-and-Language Pre-training (VLP), can enhance MMSR. In particular, we delve into two fundamental questions frequently addressed in the VLP community: (1) the architecture of VLP model [30, 28], and (2) the visual representation within VLP model [28].
6.1 Impact of Various VLP Architectures
TABLE VI: Comparison of MMSR with various VLP models of different architectures. (a): two-stream & shallow fusion. (b): single stream & deep fusion. (c): two-stream & deep fusion.
@@ 6.1 Impact of Various VLP Architectures
Our exploration initiates with a meticulous examination of diverse VLP architectures. To be specific, these can be broadly classified into three emblematic paradigms: (1) Two-stream and shallow fusion, as exemplified by models such as CLIP [26] and GroupViT [27], (2) Single-stream and deep fusion, represented by the likes of ViLT [28] and VisualBERT [29], and (3) Two-stream and deep fusion, encapsulated by models like FLAVA [30].
Specifically, we choose SASRec as the default RS architecture. For CLIP and GroupViT, we fine-tune two transformer blocks for the text and vision encoders and use a concatenation layer to connect both as the input of the downstream recommendation network. Since VisualBERT can not extract vision features itself, we use Faster-RCNN[31] with a threshold of 0.05 and select 49 regions from each image as visual input. The last 4 transformer blocks of VisualBERT are fine-tuned, which is consistent with ViLT. As for FLAVA, we fine-tune the top two transformer blocks of text and vision encoders and all fusion modules.
Results. In Table VI, we can see that MMSR performs worst in the CLIP & GroupViT architecture. By analyzing in conjunction with the other two VLPs, we hypothesize it is the shallow fusion that deteriorates the performance. We also notice that FLAVA drives MMSR to completely surpass other VLPs, which is consistent with its excellent performance in the multi-modal fields. Note, the best-performing MMSR in RQ(1) has a similar architecture to FLAVA, which may imply that the success of PTM in other domains can be transferred to the recommendation task.
6.2 Impact of Various Vision Representation
We carry out further exploration by evolving vision representation in VLPs (see Section 3 for details). We select VisualBERT [29] as our foundational model due to its iconic status within the domain of VLPs. Most subsequent strides of vision representation are indeed enhancements built upon it. To validate the MMSR, we have incorporated an assortment of vision representations, including region-based (Faster-RCNN [31]), grid-based (ResNet50 [23]), patch-Embedding (CLIP-ViT [26], excluding 12 transformers) and patch-Vision Transformer (CLIP-ViT, including 12 transformers). These representations typify a significant trajectory in the evolution of VLPs, thereby prompting our curiosity to investigate their adaptability to MMSR.
Results. As is shown in Table VII, all MMSRs with various vision representations can exceed IDSR, and patch-ViT representation is the best, which is in line with the development process in the field of VLPs. Region-based and Patch-Emd representations underperform, whereas MMSR with Grid-based and Patch-ViT demonstrate notable improvements over IDSR. This disparity might be attributed to the strong augmentation of the latter two specific vision encoders, namely ResNet50 [23] and CLIP-ViT [26]. Note that in Section 5.1, ResNet50 (grid-based) performs poorly in the previous setting. However, it achieves good results here, suggesting that deep fusion may resolve instability caused by inconsistent representations.
TABLE VII: Comparison of vision representations for MMSR.
@@ 6.2 Impact of Various Vision Representation
7 Cold-start and Transferability (Q3)
TABLE VIII: Performance comparison (%) under cold-start setting.
@@ 7 Cold-start and Transferability (Q3)
NG@10 0.1381 4.8547 4.3779 5.4046
7.1 MMSR in Cold-start Scenario
Results. Table VIII reveals that MMSR, MMSR-T, and MMSR-V all notably outperform IDSR in predicting cold items. This superior performance stems from the fact that IDSR only derives representations from behavioral patterns, but MMSR incorporates original modal features, the advantage of which becomes more apparent for cold start. Interestingly, while MMSR-T and MMSR-V have comparable performances in standard situations (as is shown in Table IV), MMSR-T tends to be more adept at handling cold items. This suggests that text-based representations might have a stronger generalizing capability for recommendation tasks than visual ones.
7.2 Results of Transfer Learning
@@
Scaling the source domain. As shown in Table IX, we choose the HM-100K, -200K, and -500K datasets as the source domain and the e-commerce Amazon-10K [55, 56] dataset as the target domain. As the scale of the source domain increases, the results of transferring increase compared with those from scratch. This implies that by increasing the sample size, MMSR can more effectively uncover user preferences within the modal content.
TABLE IX: Transfer learning from HM to Amazon. Source: HM-100K, HM-200K, HM-500K. Target: Amazon-10K Clothes Dataset. ‘Improv.’ is the relative improvement of the best transferring compared to training from scratch.
@@
NG@10 17.2302 21.7594 23.3476 22.9698 23.3558 +6.83%
Scaling the target domain. On the contrary, we select Kwai-200K as the source dataset and scale the target Bili dataset from 10K, 20K to 50K. As shown in Table X, the transferring improvement is reduced as the target domain expands.
TABLE X: Transfer learning from Kwai to Bili. Source: Kwai-200K dataset. Target: Bili-10K, -20K, and -50K dataset. ‘Improv.’ is the relative improvement of the transferring compared to the corresponding training from scratch.
@@
Results. As is shown in Table IX and Table X, MMSR can perform better than training from scratch, but it does not always hold when the ratio of the source domain to the target domain is narrowed, e.g., the results of transferring from Kwai-200K to Bili-20K and -50K show no advantages. The performance drop might stem from differences in user groups between domains. Larger downstream datasets can introduce unrelated features and noise, hampering the transfer learning performances of sequential recommenders.
Convergence. Figure 3 shows transfer learning results for MMSR, using HM-500K and Kwai-200K as sources and Amazon-10K and Bili-10K as targets. IDSR was trained for 300 epochs, while MMSR-Transfer and MMSR-Scratch only needed 60 epochs. MMSR-Transfer demonstrated superior stability, faster convergence, and better performance compared to MMSR-Scratch.
Figure 3: Comparison of MMSR-Transfer and MMSR-Scratch.
8 Bag of Tricks
@@
8.1 Full Fine-tuning vs. Top-layer Fine-tuning
In all the above experiments, we fine-tune only the top blocks for all encoders by default, which is mainly due to computational consumption. Here, we try to investigate whether all parameters need to be adapted. Table XI shows the results of full and top-layer fine-tuning showed irregular fluctuations. Among them, top-layer fine-tuning works better in MMSR. We conclude that top-layer fine-tuning is worth considering in order to balance accuracy and efficiency in practice.
TABLE XI: Full vs top-layer fine-tuning.
@@
8.2 Different Learning Rates
As shown in Table XII, using the same learning rate for all modules of the model will lead to performance degradation, which may be because the pre-trained parameters already contain a certain amount of knowledge about vision and language, and over-tuning them may lead to loss of this valuable information. As shown in Table XII, the performance of MMSR would deteriorate when all modules use the same learning rate. This is probably because the pre-trained parameters in the item encoders already contain a certain amount of knowledge about vision and language, and over-tuning may lead to the loss of this knowledge. But the downstream recommendation networks and fusion modules need training from scratch, so a large learning rate would be better.
TABLE XII: Comparison of MMSR with different learning rates for different modules. Top: TE and VE of MMSR. Downstream: the recommendation networks and fusion modules.
@@ 8.2 Different Learning Rates
8.3 Length of Texts
In our investigation, we dedicated substantial effort to analyzing the impact of varying quantities of textual information. This aspect is crucial, as the volume of textual data in a recommendation system can significantly influence its predictive accuracy. Our primary focus was on discerning how changes in the length of titles, an integral component of textual data, affected the performance of our models. As demonstrated in Table XIII, our exploration yielded noteworthy results. Both MMSR-T and MMSR exhibited a positive response to increases in title length. This trend implies that an expansion of title content can contribute to enhanced model performance. Our findings underline the significance of textual volume, particularly in title length, as a determinant factor for improving the functionality of MMSR.
TABLE XIII: The impact of various lengths of text for MMSR-T and MMSR.
@@ 8.3 Length of Texts
8.4 Mask Ratio of Images
The volume of textual and visual information can greatly affect model performance MMSR. Inspired by MAE [25], we control the image volume by adjusting its mask ratio. Table XIV shows that as the mask ratio increases, the performance of MMSR-V rises but MMSR decreases gradually. A possible reason is that information is insufficient for MMSR-V thus it would be better to keep more visual features. While for MMSR, the information may be saturated with both linguistic and visual features being fed, and it would be more intractable to optimize since the redundant image information would be taken as noise in this case.
TABLE XIV: The impact of the volume of image information for MMSR. Note that we use mask ratio from MAE [25] to control the volume of image information.